Directions: Read and answer each question carefully. For any hypothesis test, be sure to state your Ho and Ha, test statistic, p-value, and conclusion the words of the problem. For any confidence interval be sure to interpret the interval in the words of the problem. Round answers to the thousandths place.

1. A study is done to determine if Company A retains its workers longer than Company B. Company A samples 20 workers, and their average time with the company is five years with a standard deviation of 1.2 years. Company B samples 20 workers, and their average time with the company is 4.5 years with a standard deviation of 0.8. The populations are normally distributed. Is there evidence at the 5% significance level that the true mean retention time is different for A than for B? You may use a confidence interval or a hypothesis test.
Answer #1

dof=20+20-2=38

Critical Stat=t(0.975,38dof)=2.02

As t*<2.02, we fail to reject H0. Hence, there isn't sufficient evidence to say that the true mean retention time is different for than for B.

Dont forget to give a thumbs up :)